2x+3y=29
5y-4x=19

Answer by mananth(16949)

2 x + 3 y = 29 .............1
-4 x + 5 y = 19 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 3
-10 x -15 y = -145
-12 x 15 y = 57
Add the two equations
-22 x = -88
/ -22
x = 4
plug value of x in (1)
2 x + 3 y = 29
8 + 3 y = 29
3 y = 29 -8
3 y = 21
y = 7